# Terminal Velocity Matching

### v2

@@ 1 Introduction
Can we build generative models that simultaneously deliver high-quality samples, fast inference, and scalability to high-dimensional data, all from a single training stage? This is the central challenge that continues to drive research in generative models. While Diffusion Models (sohl2015deep; ho2020denoising; song2020score) and Flow Matching (liu2022flow; lipman2022flow) have become the dominant paradigms for generating images (rombach2022high; podell2023sdxl; esser2024scaling) and videos (sora; wan2025wan), they typically require many sampling steps (e.g., 50) to produce high-quality outputs. This multi-step nature makes generation computationally expensive, especially for high-dimensional data like videos.
In pursuing a single-stage training for few-step inference, recent methods have focused on directly learning integrated trajectories rather than relying on ODE solvers. Consistency-based approaches (CT (song2023consistency), CTM (kim2023consistency), sCT (lu2024simplifying)) and trajectory matching methods like MeanFlow (geng2025mean) learn to predict or match trajectory derivatives. However, these methods lack explicit connections to distribution matching, a fundamental measure of generative model quality. While Inductive Moment Matching (IMM) (zhou2025inductive) addresses this gap by providing distribution-level guarantees through Maximum Mean Discrepancy, it requires multiple particles per training step, limiting scalability.
We propose Terminal Velocity Matching (TVM), a new framework for learning ground-truth trajectories of flow-based models in a single training stage. Instead of matching time derivatives at the initial time, TVM matches them at the terminal time of trajectories. This conceptually simple shift yields powerful theoretical guarantees. We prove that our training objective upper bounds the $2$-Wasserstein distance between data and model distributions. Unlike IMM, our method provides distribution-level guarantees without requiring multiple particles. Our analysis also reveals a critical architectural limitation: current diffusion transformers (peebles2023scalable) lack Lipschitz continuity, which destabilizes TVM training. We address this with minimal architectural modifications, including RMSNorm-based QK-normalization and time embedding normalization.
@@
TVM achieves state-of-the-art results on ImageNet-$256{\mkern-1.0mu\times\mkern-1.0mu}256$, with 3.29 FID in single-step generation (outperforming MeanFlow’s (geng2025mean) with 3.43 FID) and matches/exceeds diffusion baselines with just 4 function evaluation steps (i.e., 1.99 FID for TVM vs. 2.27 FID for DiT). Similarly, our method surpasses diffusion baselines with 4-NFE on ImageNet-$512{\mkern-1.0mu\times\mkern-1.0mu}512$ (i.e. 2.94 FID for TVM vs. 3.04 FID for DiT) while outperforming prior from-scratch methods such as sCT (lu2024simplifying) and MeanFlow on single-step generation. Our method naturally interpolates between one-step and multi-step sampling without retraining, requires no training curriculum or loss modifications, and maintains stability with simple architectures. Our construction provides new insights into building scalable one/few-step generative models with distributional guarantees, demonstrating that principled theoretical design can lead to practical improvements in both training stability and generation quality.
@@ 2 Preliminaries: Flow Matching
For a given data distribution $p_{0}({\mathbf{x}}_{0})$ and prior distribution $p_{1}({\mathbf{x}}_{1})$, Flow Matching (FM) (lipman2022flow; liu2022flow) constructs a time-augmented linear interpolation ${\mathbf{x}}_{t}$ between data ${\mathbf{x}}_{0}\in\mathbb{R}^{D}$ and prior ${\mathbf{x}}_{1}\in\mathbb{R}^{D}$ such that ${\mathbf{x}}_{t}=(1-t){\mathbf{x}}_{0}+t{\mathbf{x}}_{1}$ [^2] See lipman2022flow; albergo2023stochastic for general path constructions.. For each path ${\mathbf{x}}_{t}$ conditioned on a $({\mathbf{x}}_{0},{\mathbf{x}}_{1})$ pair, there exists a conditional velocity ${\mathbf{v}}_{t}={\mathbf{x}}_{1}-{\mathbf{x}}_{0}$ for each ${\mathbf{x}}_{t}$. Under this definition, a ground-truth velocity field ${\mathbf{u}}:\mathbb{R}^{D}\times[0,1]\rightarrow\mathbb{R}^{D}$ marginal over all data and prior exists but is not known in analytical form. Therefore, a neural network ${\mathbf{u}}_{\theta}({\mathbf{x}}_{t},t)$ is used as approximation via loss
@@
For each ground-truth ${\mathbf{u}}({\mathbf{x}}_{t},t)$, there exists a corresponding displacement map $\psi:\mathbb{R}^{D}\times[0,1]\times[0,1]\rightarrow\mathbb{R}^{D}$ (i.e. flow map (boffi2024flow)) from any start time $t\in[0,1]$ to an end time $s\in[0,1]$. It is defined as the ODE integral following ${\mathbf{u}}({\mathbf{x}}_{r},r)$ for all $r\in[s,t]$, i.e.
@@ 3 Terminal Velocity Matching
where the scaling (st)(s-t) is chosen to satisfy integral boundary condition when t=st=s [^3] This is similar to CTM (kim2023consistency). See Appendix A.1 for conditions on general scaling factors.. Condition 2 can be easily enforced by FM loss (in Eq. (1)) and condition 1 can be naïvely enforced via the displacement error
@@
where notably ${\mathcal{L}}_{\text{TVM}}(\theta)$ reduces to Flow Matching objective when $t=s$ (see Appendix A.5). In practice, we employ a biased estimate of the above objective by using exponentially averaged (EMA) weights and stop-gradient for our proxy networks (li2023self). The biased per-time objective $\hat{{\mathcal{L}}}_{\text{TVM}}^{t,s}(\theta)$ is
@@
Classifier-free guidance (CFG). In the case of class-conditional generation. The ground-truth velocity field is replaced by a linear combination of class-conditional velocity ${\mathbf{u}}({\mathbf{x}}_{r},r,c)$ and unconditional velocity ${\mathbf{u}}({\mathbf{x}}_{r},r)$ (ho2022classifier), such that the new displacement map is
@@
where $w$ is the CFG weight, $c$ is class and $\varnothing$ denotes empty label. To train with CFG, we additionally condition the network on $w$ and $c$, and our class-conditional map is ${\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s,c,w)=(s-t){\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s,c,w)$ where the additional $w$ scale is chosen due to linear scaling in magnitude for marginal velocity w.r.t. $w$. The instantaneous velocity ${\mathbf{u}}_{\theta}({\mathbf{x}}_{s},s,c,w)$ is regressed against conditional velocity $w{\mathbf{v}}_{t}+(1-w){\mathbf{u}}({\mathbf{x}}_{r},r)$ where we can approximate ${\mathbf{u}}({\mathbf{x}}_{r},r)$ with our own network (chen2025visual). The per-time and per-class Flow Matching term can be modified as
@@ 4 Practical Challenges
Semi-Lipschitz control. Theorem 1 makes the crucial assumption that ${\mathbf{u}}_{\theta}({\mathbf{x}}_{s},s)$ is Lipschitz continuous. However, modern transformers with scaled dot-product attention (SDPA) and LayerNorm (LN, ba2016layer) are not Lipschitz continuous (kim2021lipschitz; qi2023lipsformer; castin2023smooth). This issue similarly applies to diffusion transformers (DiT) (peebles2023scalable). Our insight is to make minimal and non-restrictive changes to the architecture for Lipschitz control.
@@
As shown in Figure 4, the original DiT experiences training instability leading to steep jump in network activations. As a solution, we adopt RMSNorm as QK-Norm, which coinsides with the proposed ${\mathcal{L}}_{2}$ QK-Norm (qi2023lipsformer) with learnable scaling and is provably Lipschitz continuous. We also substitute all LN with RMSNorm (without learnable parameters, denoted as $\text{RMSNorm}^{-}(\cdot)$), whose Lipschitzness we show in Appendix B.1. In addition, DiT introduces Adaptive LayerNorm (AdaLN) where the output of RMSNorm is modulated by MLP outputs of time embeddings denoted as $\text{RMSNorm}^{-}(x)\odot a(t)+b(t)$ where $x$ is the input feature and $a(t),b(t)$ are scale and shift respectively. However, the Lipschitz constant of this layer depends on the magnitude of $a(t)$ which can grow unbounded and is subject to instability. We therefore employ $\text{RMSNorm}^{-}(\cdot)$ again on all modulation parameters for
@@
Figure 4 also shows the activation with our proposed changes. Activations stay smooth after our fixes. Finally, we follow qi2023lipsformer and use Lipschitz initialization for all linear layers except for time embedding layers. Note that these modifications do not explicitly constrain the Lipschitz constants of all but the key layers where instability can arise. We find such partial control of the Lipschitzness is sufficient for empirical success.
@@
where the last term involves differentiating through the network with Jacobian-Vector Product (JVP). This poses significant challenge for transformers because automatic differentiation packages, e.g. PyTorch, often do not efficiently handle JVP of SDPA. Open-source Flash Attention (dao2022flashattention) also has limited support for JVP. Crucially, different from prior works (lu2024simplifying; geng2025mean; sabour2025align), gradient is also propagated through the JVP term $\partial_{s}{\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s)$. To tackle these challenges, we propose an efficient Flash Attention kernel that (i) fuses JVP with forward pass, (ii) uses significantly less memory than naïve PyTorch attention, and (iii) supports backward pass on JVP results. We detail the implementation in Appendix C.
@@
Optimizer parameter change. Due to higher-order gradient through JVP, our loss can be subject to fluctuation with the default AdamW $\beta_{2}=0.999$. We take inspiration from language models (touvron2023llama) for mitigation and use $\beta_{2}=0.95$ to speed up update of the gradient second moment. As show in Figure 5, the terminal velocity error fluctuates significantly less after $\beta_{2}$ change.
@@
Different time distribution for FM loss. We find it empirically helpful to use a separate distribution to sample different $s$ specifically for the FM loss (see Eq. (15)), even though this may deviate from a Wasserstein interpretation. This is because we can directly transfer the proven successful time distribution from FM training for TVM. How this is used can be found in Algorithm 1 and we ablate this decision in Section 7.3.
@@ 5 Connection to Prior Works
MeanFlow. MeanFlow (geng2025mean) minimizes loss $E_{{\mathbf{x}}_{t},t,s}\left[\norm{{\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s)-F_{\text{tgt}}}_{2}^{2}\right]$ where
@@
This loss can be equivalently rewritten as $E_{{\mathbf{x}}_{t},t,s}\left[\norm{\frac{\mathrm{d}}{\mathrm{d}t}{\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s)+{\mathbf{u}}({\mathbf{x}}_{t},t)}_{2}^{2}\right]$ where ${\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s)=(s-t){\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s)$ and loss is minimized if and only if $\frac{\mathrm{d}}{\mathrm{d}t}{\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s)=-{\mathbf{u}}({\mathbf{x}}_{t},t)$ (see Appendix E.1). This exhibits duality with our proposed method in that we enforce a differential condition w.r.t. $s$ while MeanFlow differentiates w.r.t. $t$ which requires ${\mathbf{u}}({\mathbf{x}}_{t},t)$ to be propagated through JVP. In practice, ${\mathbf{u}}({\mathbf{x}}_{t},t)$ is replaced with ${\mathbf{v}}_{t}$, which introduces additional variance during training and can cause fluctuation in gradient, especially under random CFG during training (see Section 7.2). Additionally, the relationship between the loss and distribution divergence remains elusive with the introduction of ${\mathbf{v}}_{t}$. In contrast, we show our loss upper bounds $2$-Wasserstein distance up to some constant, and our theory provides the unique insight of enforcing the Lipschitzness of our network, which stablizes training.
Physics Informed Distillation (PID). PID (tee2024physics) as inspired by Physics Informed Neural Networks (raissi2019physics; cuomo2022scientific) distills pretrained diffusion models ${\mathbf{u}}_{\phi}({\mathbf{x}}_{t},t)$ into one-step samplers. It parameterizes the one-step net displacement as ${\mathbf{f}}_{\theta}({\mathbf{x}}_{1},s)=(s-1){\mathbf{u}}_{\theta}({\mathbf{x}}_{1},s)$ where ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}_{1})$ and trains via distillation loss

### v3

@@ 1 Introduction
Can we build generative models that simultaneously deliver high-quality samples, fast inference, and scalability to high-dimensional data, all from a single training stage? This is the central challenge that continues to drive research in generative models. While Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) and Flow Matching (Liu et al., 2022; Lipman et al., 2022) have become the dominant paradigms for generating images (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024) and videos (OpenAI, 2024; Wan et al., 2025), they typically require many sampling steps (e.g., 50) to produce high-quality outputs. This multi-step nature makes generation computationally expensive, especially for high-dimensional data like videos.
In pursuing a single-stage training for few-step inference, recent methods have focused on directly learning integrated trajectories rather than relying on ODE solvers. Consistency-based approaches (CT (Song et al., 2023), CTM (Kim et al., 2023), sCT (Lu and Song, 2024)) and trajectory matching methods like MeanFlow (Geng et al., 2025) learn to predict or match trajectory derivatives. However, these methods lack explicit connections to distribution matching, a fundamental measure of generative model quality. While Inductive Moment Matching (IMM) (Zhou et al., 2025) addresses this gap by providing distribution-level guarantees through Maximum Mean Discrepancy, it requires multiple particles per training step, limiting scalability.
We propose Terminal Velocity Matching (TVM), a new framework for learning ground-truth trajectories of flow-based models in a single training stage. Instead of matching time derivatives at the initial time, TVM matches them at the terminal time of trajectories. This conceptually simple shift yields powerful theoretical guarantees. We prove that our training objective upper bounds the $2$-Wasserstein distance between data and model distributions. Unlike IMM, our method provides distribution-level guarantees without requiring multiple particles. Our analysis also reveals a critical architectural limitation: current diffusion transformers (Peebles and Xie, 2023) lack Lipschitz continuity, which destabilizes TVM training. We address this with minimal architectural modifications, including RMSNorm-based QK-normalization and time embedding normalization.
@@
TVM achieves state-of-the-art results on ImageNet-$256{\mkern-1.0mu\times\mkern-1.0mu}256$, with 3.29 FID in single-step generation (outperforming MeanFlow’s (Geng et al., 2025) with 3.43 FID) and matches/exceeds diffusion baselines with just 4 function evaluation steps (i.e., 1.99 FID for TVM vs. 2.27 FID for DiT). Similarly, our method surpasses diffusion baselines with 4-NFE on ImageNet-$512{\mkern-1.0mu\times\mkern-1.0mu}512$ (i.e. 2.94 FID for TVM vs. 3.04 FID for DiT) while outperforming prior from-scratch methods such as sCT (Lu and Song, 2024) and MeanFlow on single-step generation. Our method naturally interpolates between one-step and multi-step sampling without retraining, requires no training curriculum or loss modifications, and maintains stability with simple architectures. Our construction provides new insights into building scalable one/few-step generative models with distributional guarantees, demonstrating that principled theoretical design can lead to practical improvements in both training stability and generation quality.
@@ 2 Preliminaries: Flow Matching
For a given data distribution $p_{0}({\mathbf{x}}_{0})$ and prior distribution $p_{1}({\mathbf{x}}_{1})$, Flow Matching (FM) (Lipman et al., 2022; Liu et al., 2022) constructs a time-augmented linear interpolation ${\mathbf{x}}_{t}$ between data ${\mathbf{x}}_{0}\in\mathbb{R}^{D}$ and prior ${\mathbf{x}}_{1}\in\mathbb{R}^{D}$ such that ${\mathbf{x}}_{t}=(1-t){\mathbf{x}}_{0}+t{\mathbf{x}}_{1}$ [^2] See Lipman et al. (2022); Albergo et al. (2023) for general path constructions.. For each path ${\mathbf{x}}_{t}$ conditioned on a $({\mathbf{x}}_{0},{\mathbf{x}}_{1})$ pair, there exists a conditional velocity ${\mathbf{v}}_{t}={\mathbf{x}}_{1}-{\mathbf{x}}_{0}$ for each ${\mathbf{x}}_{t}$. Under this definition, a ground-truth velocity field ${\mathbf{u}}:\mathbb{R}^{D}\times[0,1]\rightarrow\mathbb{R}^{D}$ marginal over all data and prior exists but is not known in analytical form. Therefore, a neural network ${\mathbf{u}}_{\theta}({\mathbf{x}}_{t},t)$ is used as approximation via loss
@@
For each ground-truth ${\mathbf{u}}({\mathbf{x}}_{t},t)$, there exists a corresponding displacement map $\psi:\mathbb{R}^{D}\times[0,1]\times[0,1]\rightarrow\mathbb{R}^{D}$ (i.e. flow map (Boffi et al., 2024)) from any start time $t\in[0,1]$ to an end time $s\in[0,1]$. It is defined as the ODE integral following ${\mathbf{u}}({\mathbf{x}}_{r},r)$ for all $r\in[s,t]$, i.e.
@@ 3 Terminal Velocity Matching
where the scaling (st)(s-t) is chosen to satisfy integral boundary condition when t=st=s [^3] This is similar to CTM (Kim et al., 2023). See Appendix A.1 for conditions on general scaling factors.. Condition 2 can be easily enforced by FM loss (in Eq. (1)) and condition 1 can be naïvely enforced via the displacement error
@@
where notably ${\mathcal{L}}_{\text{TVM}}(\theta)$ reduces to Flow Matching objective when $t=s$ (see Appendix A.5). In practice, we employ a biased estimate of the above objective by using exponentially averaged (EMA) weights and stop-gradient for our proxy networks (Li et al., 2023). The biased per-time objective $\hat{{\mathcal{L}}}_{\text{TVM}}^{t,s}(\theta)$ is
@@
Classifier-free guidance (CFG). In the case of class-conditional generation. The ground-truth velocity field is replaced by a linear combination of class-conditional velocity ${\mathbf{u}}({\mathbf{x}}_{r},r,c)$ and unconditional velocity ${\mathbf{u}}({\mathbf{x}}_{r},r)$ (Ho and Salimans, 2022), such that the new displacement map is
@@
where $w$ is the CFG weight, $c$ is class and $\varnothing$ denotes empty label. To train with CFG, we additionally condition the network on $w$ and $c$, and our class-conditional map is ${\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s,c,w)=(s-t){\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s,c,w)$ where the additional $w$ scale is chosen due to linear scaling in magnitude for marginal velocity w.r.t. $w$. The instantaneous velocity ${\mathbf{u}}_{\theta}({\mathbf{x}}_{s},s,c,w)$ is regressed against conditional velocity $w{\mathbf{v}}_{t}+(1-w){\mathbf{u}}({\mathbf{x}}_{r},r)$ where we can approximate ${\mathbf{u}}({\mathbf{x}}_{r},r)$ with our own network (Chen et al., 2025). The per-time and per-class Flow Matching term can be modified as
@@ 4 Practical Challenges
Semi-Lipschitz control. Theorem 1 makes the crucial assumption that ${\mathbf{u}}_{\theta}({\mathbf{x}}_{s},s)$ is Lipschitz continuous. However, modern transformers with scaled dot-product attention (SDPA) and LayerNorm (LN, Ba et al. (2016)) are not Lipschitz continuous (Kim et al., 2021; Qi et al., 2023; Castin et al., 2023). This issue similarly applies to diffusion transformers (DiT) (Peebles and Xie, 2023). Our insight is to make minimal and non-restrictive changes to the architecture for Lipschitz control.
@@
As shown in Figure 4, the original DiT experiences training instability leading to steep jump in network activations. As a solution, we adopt RMSNorm as QK-Norm, which coinsides with the proposed ${\mathcal{L}}_{2}$ QK-Norm (Qi et al., 2023) with learnable scaling and is provably Lipschitz continuous. We also substitute all LN with RMSNorm (without learnable parameters, denoted as $\text{RMSNorm}^{-}(\cdot)$), whose Lipschitzness we show in Appendix B.1. In addition, DiT introduces Adaptive LayerNorm (AdaLN) where the output of RMSNorm is modulated by MLP outputs of time embeddings denoted as $\text{RMSNorm}^{-}(x)\odot a(t)+b(t)$ where $x$ is the input feature and $a(t),b(t)$ are scale and shift respectively. However, the Lipschitz constant of this layer depends on the magnitude of $a(t)$ which can grow unbounded and is subject to instability. We therefore employ $\text{RMSNorm}^{-}(\cdot)$ again on all modulation parameters for
@@
Figure 4 also shows the activation with our proposed changes. Activations stay smooth after our fixes. Finally, we follow Qi et al. (2023) and use Lipschitz initialization for all linear layers except for time embedding layers. Note that these modifications do not explicitly constrain the Lipschitz constants of all but the key layers where instability can arise. We find such partial control of the Lipschitzness is sufficient for empirical success.
@@
where the last term involves differentiating through the network with Jacobian-Vector Product (JVP). This poses significant challenge for transformers because automatic differentiation packages, e.g. PyTorch, often do not efficiently handle JVP of SDPA. Open-source Flash Attention (Dao et al., 2022) also has limited support for JVP. Crucially, different from prior works (Lu and Song, 2024; Geng et al., 2025; Sabour et al., 2025), gradient is also propagated through the JVP term $\partial_{s}{\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s)$. To tackle these challenges, we propose an efficient Flash Attention kernel that (i) fuses JVP with forward pass, (ii) uses significantly less memory than naïve PyTorch attention, and (iii) supports backward pass on JVP results. We detail the implementation in Appendix C.
@@
Optimizer parameter change. Due to higher-order gradient through JVP, our loss can be subject to fluctuation with the default AdamW $\beta_{2}=0.999$. We take inspiration from language models (Touvron et al., 2023) for mitigation and use $\beta_{2}=0.95$ to speed up update of the gradient second moment. As show in Figure 5, the terminal velocity error fluctuates significantly less after $\beta_{2}$ change.
@@
Different time distribution for FM loss. We find it empirically helpful to use a separate distribution to sample different $s$ specifically for the FM loss (see Eq. (15))This is because we can directly transfer the proven successful time distribution from FM training for TVM. How this is used can be found in Algorithm 1 and we ablate this decision in Section 7.3.
@@ 5 Connection to Prior Works
MeanFlow. MeanFlow (Geng et al., 2025) minimizes loss $E_{{\mathbf{x}}_{t},t,s}\left[\norm{{\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s)-F_{\text{tgt}}}_{2}^{2}\right]$ where
@@
This loss can be equivalently rewritten as $E_{{\mathbf{x}}_{t},t,s}\left[\norm{\frac{\mathrm{d}}{\mathrm{d}t}{\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s)+{\mathbf{u}}({\mathbf{x}}_{t},t)}_{2}^{2}\right]$ where ${\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s)=(s-t){\mathbf{F}}_{\theta}({\mathbf{x}}_{t},t,s)$ and loss is minimized if and only if $\frac{\mathrm{d}}{\mathrm{d}t}{\mathbf{f}}_{\theta}({\mathbf{x}}_{t},t,s)=-{\mathbf{u}}({\mathbf{x}}_{t},t)$ (see Appendix E.1). This exhibits duality with our proposed method in that we enforce a differential condition w.r.t. $s$ while MeanFlow differentiates w.r.t. $t$ which requires ${\mathbf{u}}({\mathbf{x}}_{t},t)$ to be propagated through JVP. In practice, ${\mathbf{u}}({\mathbf{x}}_{t},t)$ is replaced with ${\mathbf{v}}_{t}$, which introduces additional variance during training and can cause fluctuation in gradient, especially under random CFG during training (see Section 7.2). Additionally, the relationship between the loss and distribution divergence remains elusive with the introduction of ${\mathbf{v}}_{t}$. In contrast, we show our loss upper bounds $2$-Wasserstein distance, which provides unique insights on smoothness properties of our network.
Physics Informed Distillation (PID). PID (Tee et al., 2024) as inspired by Physics Informed Neural Networks (Raissi et al., 2019; Cuomo et al., 2022) distills pretrained diffusion models ${\mathbf{u}}_{\phi}({\mathbf{x}}_{t},t)$ into one-step samplers. It parameterizes the one-step net displacement as ${\mathbf{f}}_{\theta}({\mathbf{x}}_{1},s)=(s-1){\mathbf{u}}_{\theta}({\mathbf{x}}_{1},s)$ where ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}_{1})$ and trains via distillation loss